Calculate : 5x - 1/2 if 2x + 3 = 19
Answer by checkley77(12844)

2x+3=19
2x=19-3
2x=16
x=16/2
x=8 then:
5x-1/2
5*8-1/2
40-1/2
39 1/2 answer